Fermi National Laboratory
 

Accelerator Update for November 7, 2007 - November 9, 2007

 

Wednesday November 7

The day shift began with the Tevatron (TeV) in access for the sector B0 low beta magnet moves, with the Recycler preparing to stash, with the Antiproton Source (Pbar) stacking (22.5E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access, and with beam to the SciBooNE / MiniBooNE experiments.

At 8:26 AM, Operations reported that Pbar was having a stacking efficiency problem, but it turned out to be a DCCT signal problem.

At 9:13 AM, Sci/MiniBooNE tripped off. It reset.

At 10:05 AM, the 750 KeV area buncher tripped off. It reset.

Operations reported at 10:47 AM that interlock technicians were accessing the Meson M03&4 enclosures to check on ODH equipment.

A MI expert began a multibatch slipstacking study at 1:30 PM.

At 2:18 PM, the Booster West Anode Power Supply (WAPS) tripped off. It reset.

At 2:22 PM, Sci/MiniBooNE tripped off. It reset.

The MI study ended at 3:13 PM.

Operations reported at 4 PM that the Recycler had stashed antiprotons once during the day shift.

At 4:02 PM, Sci/MiniBooNE tripped off. It reset. Operators reported that every time a certain event went out on the Time Line Generator (TLG) a power supply (E:MBEXPS) tripped off. Experts reported that they had added a filter to clean up the signal earlier in the day, but hadn’t changed its voltage limits in the permit system.

A TeV expert reported at 4:45 PM that the TeV sector A side of the low beta magnets at CDF had been aligned. He said that it would take an hour or so to reinstall the shielding and then begin the cryo system cooldown.

At 4:48 PM, Operations reported that WAPS had tripped off. Experts investigated. They had it back on line about 20 minutes later.

A TeV expert reported at 7:04 PM that the TeV sector B alignment of the low beta magnets and replacing its shielding had all been completed. Its cryo system cool down was just starting.

At 7:27 PM, Sci/MiniBooNE tripped off for the first of two times in 90 minutes. Operators contacted an expert after both trips and then adjusted its limits.

A Pbar expert began a reverse proton study at 7:51 PM.

At 8:13 PM, a TeV vacuum pump (T:F1RP3) tripped off. Operations contacted vacuum experts.

Operations reported at 8:30 PM that Operators had completed their search and secure of the CDF and D0 collision halls.

Operations reported at 10:40 PM that Operators had completed their search and secure of the TeV tunnels.

At 11:54 PM, Booster suffered a vacuum burst that closed vacuum valves and tripped off a power supply (GMPS) and the Booster chopper. Everything reset.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Thursday November 8

The midnight shift began with Operations monitoring the stash (65.5E10) and the stack (20.2E10), and with Sci /MiniBooNE taking beam.

Pbar completed their study and resumed stacking at 12:43 AM.

Operations put the TeV through a dry squeeze at 3:37 AM.

TeV experts began a study period at 4:35 AM.

At 7:21 AM, the TeV suffered a low beta quench at sector D1. Experts investigated.

Operations reported at 8:03 AM that the Recycler had stashed twice during the midnight shift.

Pbar experts began a study at 8:18 AM.

The Cryo Coordinator reported at 8:57 AM that the TeV sector B1 Thermometry crate had noisy readbacks. An expert investigated and made adjustments.

SwitchYard experts began trying to establish beam to the M01 absorber around 9:45 AM.

At 9:48 AM, MI aborted due to a LCW temperature problem at MI-30. An expert investigated.

At 10:46 the 750 KeV area buncher tripped off. It reset.

The Neutron Therapy Facility reported at 10:54 AM that they had treated their last patient for the day.

Operations reported at 11:15 AM that they had received permission to send beam to MTest.

Operations established beam to MTest at 11:24 AM.

At 11:27 AM, the Sci/MiniBooNE tripped off due to MBEXPS. An expert investigated and had beam back on about an hour later.

Operations reported at 12:30 PM that a Recycler expert was working on their damper system.

Operations ramped the TeV to 980 GeV at 1:28 PM.

Operations put the TeV into a dry squeeze at 1:40 PM.

A MI expert began a slipstacking study at 1:46 PM.

The Pbar study ended and stacking resumed at 1:56 PM.

TeV experts began a study at 2:23 PM.

At 3:33 PM, Operations reported that a MI RF station (MIRF10) had tripped off three times with no indication of a problem. Operations contacted an expert.

Operations reported at 4 PM that the Recycler had stashed once during the day shift.

At 4:22 PM, the Linac low energy buncher tripped off. It reset.

At 4:54 PM, Operations put Pbar into standby while two RF stations were repaired. The work was completed by 5:28 PM.

Pbar experts resumed their study at 6:13 PM.

Operations put the TeV through a dry squeeze at 6:27 PM.

MTest experimenters reported at 8:29 PM that they were done with the beam for the night.

At 9:31 PM, the TeV quenched at sector F1. Experts investigated and later reported that the quench was due to some changes made to a sequencer earlier in the day.

Operations reported at 11:58 PM that the Recycler had stashed once during the evening shift.

Friday November 9

The midnight shift began with Operations monitoring the stash (193.3E10) and stack (128.7E10), and with Sci /MiniBooNE taking beam.

At 2:06 AM, the TeV quenched at sector F1. Experts reported that there were high voltages on F13U and F17U due to a VFC chassis problem at F1. Operators replaced the chassis.

At 4:08 AM, Operations reported that there was still a problem with the VFC chassis at sector F1. They called in an expert.

At 5:19 AM, a CUB inlet valve closed, which caused the TeV CUB outlet valve to close.

At 5:37 AM, the Linac beamstop closed.

At 5:38 AM, Operations reported that the TeV LCW valves closed at F17.

At 5:40 AM, Operations reported that all these valve closings were due to a loss of nitrogen pressure. The nitrogen DEWAR at TeV sector A0 had emptied. An expert reported that there was supposed to have been a delivery yesterday.

At 5:49 AM, the 750 KeV area beamstop closed.

At 6:05 AM, Operations turned the Linac high and low level RF systems to zero due to the LCW problems. Later, Operations turned off the Linac quadrupole magnets and power supplies.

At 6:16 AM, Operations put Pbar into standby.

At 6:23 AM, Operations had the TeV LCW pumps turned off.

At 6:26 AM, Operations reported that Booster was off.

At 6:29 AM, Operations reported that the P1 & 2 transfer lines, as well as the Pbar AP1 line, were turned off.

At 6:49 AM, an expert reported that he had talked to the nitrogen vendor who said they had an order to deliver only during daylight hours. The expert stated that this wasn’t in the contract and had never been a problem before. The Cryo Coordinator reported that he would direct their next nitrogen delivery, due around 8:30 AM, to A0.

At 7:32 AM, Operations reported that 80% of their vacuum pumps had tripped off.

Operations prepared the Booster and the Pelletron for access around 7:52 AM.

The Plans for Friday and the Weekend

The plans for today are to recover from loss of nitrogen pressure and the resulting LCW and vacuum problems. The TeV will conduct a couple of shifts of studies. The next store will go into the TeV on Saturday.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/09/2007 -->    email Fermilab